OpenStudy (anonymous):

Solve x3 = 64 over 27.

OpenStudy (anonymous):

@master_splinter

OpenStudy (anonymous):

x =(-12-√-432)/18=(-2-2i√ 3 )/3= -0.6667-1.1547i x =(-12+√-432)/18=(-2+2i√ 3 )/3= -0.6667+1.1547i x = 4/3

OpenStudy (anonymous):

Thanks, could you explain that please, i don't know if that answer would be correct, but thanks for the help.

OpenStudy (anonymous):

We can move the 3 across by dividing each side by 3, then we're left with x = 64 over (27 times 3) which again equals 64 over 81

OpenStudy (anonymous):

The answer would be 4 over 3

OpenStudy (anonymous):

Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation :

OpenStudy (anonymous):

Can anyone help me with two more, i will open a new thread.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

sure thing

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

If you're going to write x in the power of 3, you do x^3. x3 or 3x means 3 times x, it's not the same.